OpenStudy (anonymous):

Simplify the following expression. 1 1/2*2 1/6 A. 3 1/4 B. 1 8/11 C. 19/30 D. 8 2/11 @imqwerty @itsbrobro

OpenStudy (anonymous):

@itsbribro

OpenStudy (anonymous):

@jabez177

OpenStudy (alexandervonhumboldt2):

hey this is the earh science section you posted this in wrong section but anyways i gonna help you

jabez177 (jabez177):

This goes in the math Section.

OpenStudy (alexandervonhumboldt2):

first convert to imporper fraction

OpenStudy (anonymous):

oh sorry :)

OpenStudy (alexandervonhumboldt2):

\[a \frac{ b }{ c }=\frac{ a*c+b }{ c }\] \[1\frac{ 1 }{ 2 }=\frac{ 3 }{ 2 }\] \[2\frac{ 1 }{ 6 }=\frac{ 13 }{ 6 }\]

OpenStudy (anonymous):

do i make the denominators the same?

OpenStudy (alexandervonhumboldt2):

now multply \[\frac{ 3 }{ 2 }*\frac{ 13 }{ 6 }=\frac{ 3*13 }{ 6*2 }\]

OpenStudy (alexandervonhumboldt2):

no, you need to convert to improper fraction

OpenStudy (anonymous):

so it make 39/18

OpenStudy (anonymous):

i mean 39 /12

OpenStudy (alexandervonhumboldt2):

yes but this can be simpllfied: \[\frac{ 39 }{ 12 }=\frac{13 }{ 4 }\]

OpenStudy (alexandervonhumboldt2):

you could have just canceled the 3

OpenStudy (anonymous):

okay so ........?